# ControlFusion: A Controllable Image Fusion Framework with Language-Vision Degradation Prompts

### v1

@@ Abstract
Current image fusion methods struggle to address the composite degradations encountered in real-world imaging scenarios and lack the flexibility to accommodate user-specific requirements. In response to these challenges, we propose a controllable image fusion framework with language-vision prompts, termed ControlFusion, which adaptively neutralizes composite degradations. On the one hand, we develop a degraded imaging model that integrates physical imaging mechanisms, including the Retinex theory and atmospheric scattering principle, to simulate composite degradations, thereby providing potential for addressing real-world complex degradations from the data level. On the other hand, we devise a prompt-modulated restoration and fusion network that dynamically enhances features with degradation prompts, enabling our method to accommodate composite degradation of varying levels. Specifically, considering individual variations in quality perception of users, we incorporate a text encoder to embed user-specified degradation types and severity levels as degradation prompts. We also design a spatial-frequency collaborative visual adapter that autonomously perceives degradations in source images, thus eliminating the complete dependence on user instructions. Extensive experiments demonstrate that ControlFusion outperforms SOTA fusion methods in fusion quality and degradation handling, particularly in countering real-world and compound degradations with various levels.
Figure 1: Fusion performance comparison across real-world, composite degradation, and varying degradation levels.
@@ 1 Introduction
Image fusion is a crucial technique in image processing. By effectively leveraging complementary information from multiple sources, the limitations associated with information collection by single-modal sensors can be significantly mitigated [40]. Among the various research areas within this field, infrared-visible image fusion (IVIF) has garnered considerable attention. IVIF integrates essential thermal information from infrared (IR) images with the intricate texture details from visible (VI) images, offering a richer and more complete depiction of the scene [44]. By harmonizing diverse information and producing visually striking results, IVIF has found extensive applications in areas such as security surveillance [45], military detection [23], scene understanding [7, 43], and assisted driving [1]etc.
@@
Recently, IVIF has emerged as a focal point of research, leading to rapid advancements in related methods. Based on the architecture of deep networks, these methods can be categorized into convolutional neural network- [14, 50], autoencoder- [11, 12], generative adversarial network- [19, 15], Transformer- [21, 42], and diffusion model-based [49, 41] methods. Besides, from a functional perspective, they can be categorized into visual-oriented [30, 50], joint registration-fusion [28, 37], semantic-driven [29, 16], and degradation-robust [32, 41] schemes.
@@
Notably, although degradation-robust schemes can mitigate degradation interference to some extent, they still have certain limitations. On the one hand, these methods often rely on simplistic approaches to construct training data, overlooking the domain gap between simulated data and realistic images, which significantly hampers their generalizability in practical scenarios, as presented in Fig. 1(I). On the other hand, they are tailored for specific or single types of degradation, making them ineffective in handling more complex composite degradations, as illustrated in Fig. 1(II). Last but not least, as shown in Fig. 1(III), existing methods lack degradation level modeling, causing a sharp decline in performance as degradation intensifies. Furthermore, these approaches do not offer the flexibility to dynamically adjust fusion results according to the diverse perceptual requirements of different users.
@@
To overcome these limitations, we propose a versatile and controllable fusion model based on language-vision prompts, termed ControlFusion. On the one hand, we introduce a physics-driven degraded imaging model, which accurately simulates the physical mechanisms of the actual degradation process, e.g., illumination-, weather-, and sensor-related distortions. The proposed model not only effectively bridges the gap between synthetic data and real-world images but also offers valuable data support for handling complex composite degradations. On the other hand, we develop a prompt-modulated image restoration and fusion network to generate high-quality fusion results. The prompt-modulated module enables our network to dynamically adjust feature distribution based on degradation characteristics (which can be specified by users), achieving robust feature enhancement. This also allows our method to respond to diverse user requirements. Furthermore, we devise a spatial-frequency visual adapter that combines frequency-domain degradation priors to directly extract degradation cues from degraded inputs, eliminating the heavy reliance on user instructions, as customizing prompts for each scene is time-consuming and labor-intensive. As shown in Fig. 1, benefiting from the aforementioned designs, our method excels in handling both real-world and composite degradation scenarios, and it shows strong robustness across various degradation levels. In summary, our main contributions are as follows:
@@
    We propose ControlFusion, a versatile image restoration and fusion framework that uniformly models diverse degradation types and degrees, using textual and visual prompts as a medium. Specifically, its controllability enables it to respond to user-specific customization needs.
@@
    A spatial-frequency visual adapter is devised to integrate frequency characteristics and directly extract text-aligned degradation prompts from visual images, enabling large-scale automated deployment.
@@
    A physics-driven imaging model is developed, integrating physical mechanisms such as the Retinex theory and atmospheric scattering principle to bridge the gap between synthetic data and real-world images, while providing data support for addressing composite degradations.
@@ 2 Related Work
Image Fusion. Earlier visual-oriented fusion approaches primarily concentrated on merging complementary information from multiple modalities and improving visual fidelity. These methods often depend on sophisticated network architectures and carefully designed loss functions to ensure that the integrated content retains the essential characteristics of the original images, preserving their unique features while enhancing overall quality. The mainstream network architectures primarily include CNN-based [20, 47], AE-based [12, 48], GAN-based [19, 15], Transformers [21] and diffusion models [49]. Furthermore, several schemes including joint registration and fusion [33, 37], semantic-driven [29, 15], and degradation-robust [32, 38, 41] methods, are proposed to broaden the practical applications of image fusion. For instance, Tang et al. [31] and Liu et al. [17] developed corresponding solutions for illumination distortions and noise interference, respectively. In addition, Yi et al. [38] proposed a unified image restoration and fusion network to address various degradations. However, they are unable to handle mixed degradations and struggle to generalize to real-world scenarios. In particular, Text-IF [38] relies on tedious manual efforts to customize text prompts for each scene, hindering large-scale automated deployment.
Image Restoration with Vision-Language Models. With the advancement of deep learning toward multimodal integration, the image restoration field has progressively evolved into a new paradigm of text-driven schemes. CLIP [26] establishes visual-textual semantic representation alignment through dual-stream Transformer encoders pre-trained on 400 million image-text pairs. This framework lays the foundation for the widespread adoption of Prompt Engineering in computer vision. Recent studies integrate CLIP encoders with textual user instructions, proposing generalized restoration frameworks including PromptIR [24], AutoDIR [9], and InstructIR [2], which effectively handle diverse degradation types. Moreover, SPIRE [25] further introduces fine-grained textual restoration cues by quantifying degradation severity levels and supplementing semantic information for precision restoration. To eliminate reliance on manual guidance, Luo et al. [18] developed DA-CLIP by fine-tuning CLIP on mixed-degradation datasets, enabling degradation-aware embeddings to facilitate distortion handling. However, existing unified restoration methods are primarily designed for natural images, which exhibit notable limitations when processing multimodal image pairs (e.g., infrared-visible).
Figure 2: The overall framework of our semantic and degradation dual-prior guided image fusion network.
@@ 3 Physics-driven Degraded Imaging Model
3.1 Imaging Formulation
In real-world scenarios, images are often affected by multiple degradation factors simultaneously, which leads to suboptimal performance of restoration and fusion methods designed for isolated degradations. To tackle the complex and variable degradation challenges, we propose a physics-driven imaging model aimed at reducing the domain gap between simulated data and real-world imagery. Visible (VI) images are typically degraded by illumination conditions (low-light, overexposure), weather (rain, haze), and sensor-related issues (noise, blur), while infrared (IR) images usually suffer from sensor-related interference, such as stripe noise and low contrast. Thereby, given a clear image $I_{m}(m\in\{ir,vi\})$, the proposed imaging model can be mathematically represented as:
@@ 3.1 Imaging Formulation
3.2 Composite Degradation Dataset with Levels
Based on Eq. (1), we construct a multi-modal composite Degradation Dataset with four Levels (DDL-12), comprising 12 distinct degradations along with their corresponding high-quality images. In particular, we selecte $2,050$ high-quality clear images from multiple publicly available infrared-visible datasets [36, 16, 8, 30]. Subsequently, the degraded imaging model is employed to synthesize degraded images with 12 types and 4 levels of degradation. Finally, the DDL-12 dataset contains approximately $32,000$ training image pairs and $4,000$ test image pairs.
@@ 4.1 Problem Formulation
Given two source images $I_{ir}\in\mathbb{R}^{H\times W\times 1}$ and $I_{vi}\in\mathbb{R}^{H\times W\times 3}$, the typical fusion paradigm employs a deep network $\mathcal{N}_{f}$ to integrate complementary information and synthesize the fused image $I_{f}$, expressed as:
$I_{f}=\mathcal{N}_{f}(I_{ir},I_{vi}).$ (6)
However, in complex scenarios, source images usually suffer from degradation interference, thus the advanced fusion paradigm must take both image restoration and information aggregation into account. Concatenation and coupling are two common approaches to implement this. Notably, while the concatenation approach is straightforward, it neglects the intrinsic relationship between restoration and fusion, frequently yielding suboptimal outcomes. As illustrated in Fig. 2, we propose a controllable paradigm that couples restoration and fusion with degradation prompt, termed ControlFusion, to overcome this limitation. On the one hand, the coupled approach enhances task synergy. On the other hand, leveraging degradation prompts to modulate the restoration and fusion process enables the network to adapt to diverse degradation distributions while meeting user-specific customization requirements. Thus, the proposed restoration and fusion paradigm can be defined as:
$I_{f}=\mathcal{N}_{rf}(I_{ir},I_{vi},p~{}|~{}\Omega),$ (7)
where $\mathcal{N}_{rf}$ is a restoration and fusion network, $P$ denotes degradation prompts, and $\Omega$ indicates the degradation set. Notably, degradation prompts can be either user-specified text or directly extracted from input images, greatly enhancing the flexibility of our framework without heavily relying on user instructions. To fully harness its potential, our ControlFusion employs a two-stage optimization protocol. Stage I aligns textual prompts with visual embedding, while Stage II optimizes the overall framework.
@@ 4.2 Stage I: Textual-Visual Prompts Alignment
Text Semantic Encoder. Adopting text to represent the degradation type and degree is the most straightforward approach. Thus, we introduce a pre-trained vision-language model (i.e., CLIP [26]), to achieve prompt embedding. Given a user instruction $T_{instr}$, the text semantic encoder $\mathcal{E}_{t}$ from CLIP is deployed to convert it into a text embedding:
$p_{text}=\mathcal{E}_{t}(T_{instr}).$ (8)
@@
Spatial-Frequency Collaborative Visual Adapter. It is important to emphasize that customizing text instructions for each image is extremely time-consuming and labor-intensive. Therefore, we expect to develop a textual-visual adapter that directly extracts visual embeddings from source images that align with $p_{text}$, enabling automated deployment. In particular, as shown in Fig. 2, images with different degradations in the same scene exhibit distinct spectral characteristics, indicating that frequency features contain rich degradation priors. To this end, we devise a spatial-frequency collaborative visual adapter (SFVA) to extract visual prompts. In the frequency branch, we utilize the Fast Fourier Transform (FFT) algorithm in conjunction with a CNN to extract frequency features $F_{fre}$:
$F_{fre}^{m}$ $=\textstyle{\sum_{x=0}^{W-1}\sum_{y=0}^{H-1}D_{m}(x,y)e^{-j2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)},}$ (9)
$F_{fre}$ $=\text{Linear}\left(\text{Conv}\left([F_{fre}^{ir},F_{fre}^{vi}]\right)\right),$
where [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes channel-wise concatenation. In the spatial branch, we employ cropping and downsampling for data augmentation, followed by a CNN-based visual encoder to extract spatial features $F_{spa}$. Then, we concatenate $F_{fre}$ and $F_{spa}$, and apply successive linear layers to obtain the final visual embedding $p_{vis}$. Furthermore, we employ MSE loss $\mathcal{L}_{mse}$ and cosine similarity loss $\mathcal{L}_{cos}$ to impose constraints on the similarity between $p_{vis}$ and the pre-defined $p_{text}$. The loss for Stage I is formulated as:
where $\lambda_{1}$ and $\lambda_{2}$ are parameters, balancing various losses.
@@ 4.3.1 Network Architectures
Feature Encoding and Fusion Layer. As shown  Fig. 2, we devise the hierarchical transformer-based encoders to extract multi-scale feature representations from degraded infrared ($D_{ir}$) and visible ($D_{vi}$) images separately, which is formulated as:
$\{F_{ir},F_{vi}\}=\{\mathcal{E}_{ir}(D_{ir}),\mathcal{E}_{vi}(D_{vi})\},$ (11)
where $\mathcal{E}_{ir}$ and $\mathcal{E}_{vi}$ are infrared and visible image encoders.
@@
$F_{f}^{ir}\!=\!\text{softmax}\Big{(}\!\frac{Q_{vi}K_{ir}}{\sqrt{d_{k}}}\!\Big{)}V_{ir},F_{f}^{vi}\!=\!\text{softmax}\Big{(}\!\frac{Q_{vi}K_{ir}}{\sqrt{d_{k}}}\!\Big{)}V_{vi},$ (13)
where $d_{k}$ is a learnable scaling factor. Finally, we concatenate $F_{f}^{ir}$ and $F_{f}^{vi}$ to get fusion features: $F_{f}=[F_{f}^{ir},F_{f}^{vi}]$.
Prompt-modulated Module. To dynamically adapt fusion feature distributions to varying degradation patterns and degrees, we propose a prompt-modulated module (PMM) for robust feature enhancement. First, sequential MLPs ($\Phi_{p}$) are utilized to derive distributional optimization parameters:
$[\gamma_{p},\beta_{p}]=\Phi_{p}(p),\quad\text{where}\ p\in\{p_{\text{vis}},p_{\text{text}}\}.$
Then, $\gamma_{p}$ and $\beta_{p}$ are applied for feature scaling and bias shifting in a residual manner:
$\hat{F}_{f}=(1+\gamma_{p})\odot F_{f}+\beta_{p},$ (14)
where $\odot$ denotes the Hadamard product, and $\hat{F}_{f}$ indicates the enhanced features incorporating degradation prompts. Image Decoder. We further deploy a series of Transformer-based decoder $\mathcal{D}_{f}$ with the self-attention mechanism to progressively reconstruct the fused image $I_{f}$:
$I_{f}=\mathcal{D}(\hat{F}_{f}).$ (15)
In particular, PMM and $\mathcal{D}_{f}$ are tightly coupled in a multi-stage process, effectively incorporating fusion features and degradation prompts to synthesize the desired image.
MSRS LLVIP RoadScene FMB
Methods EN SD VIF Qabf EN SD VIF Qabf EN SD VIF Qabf EN SD VIF Qabf
DDFM [49] 6.431 47.815 0.844 0.643 6.914 48.556 0.693 0.517 6.994 47.094 0.775 0.595 6.426 40.597 0.495 0.442
DRMF [32] 6.268 45.117 0.669 0.550 6.901 50.736 0.786 0.626 6.231 44.221 0.728 0.527 6.842 41.816 0.578 0.372
EMMA [50] 6.747 52.753 0.886 0.605 6.366 47.065 0.743 0.547 6.959 46.749 0.698 0.664 6.788 38.174 0.542 0.436
LRRNet [12] 6.761 49.574 0.713 0.667 6.191 48.336 0.864 0.575 7.185 46.400 0.756 0.658 6.432 48.154 0.501 0.368
SegMiF [16] 7.006 57.073 0.764 0.586 7.260 45.892 0.539 0.459 6.736 48.975 0.629 0.584 6.363 47.398 0.539 0.482
Text-IF [38] 6.619 55.881 0.753 0.656 6.364 49.868 0.859 0.566 6.836 47.596 0.634 0.609 7.397 47.726 0.568 0.528
Text-Difuse [41] 6.990 56.698 0.850 0.603 7.546 55.725 0.883 0.659 6.826 50.230 0.683 0.662 6.888 49.558 0.793 0.653
Table 2: Quantitative comparison results in different degraded scenarios with pre-enhancement.
VI (Blur) VI (Rain) VI (Low-light, LL) VI (Over-exposure, OE)
DDFM [49] 0.141 39.421 39.047 35.411 0.191 38.836 46.285 36.376 0.156 39.495 41.782 31.759 0.143 43.167 43.440 32.099
DRMF [32] 0.128 40.739 40.968 40.722 0.174 48.164 48.565 41.174 0.143 41.428 37.947 38.287 0.190 48.334 42.582 44.256
EMMA [50] 0.131 43.472 41.744 42.553 0.138 45.824 44.916 43.378 0.158 39.674 44.827 40.857 0.180 46.731 47.616 40.242
LRRNet [12] 0.163 42.981 37.268 45.389 0.185 43.291 41.891 46.285 0.164 40.486 34.836 41.639 0.160 42.548 48.414 42.190
SegMiF [16] 0.152 43.005 43.516 44.000 0.195 40.528 49.094 44.274 0.177 44.073 48.376 44.829 0.166 49.132 38.019 38.484
Text-IF [38] 0.164 44.801 46.542 48.401 0.164 41.287 47.380 49.298 0.163 41.096 49.174 47.257 0.172 40.298 45.599 47.330
Text-Difuse [41] 0.172 44.958 47.699 46.376 0.173 39.243 50.017 47.297 0.192 46.734 50.126 49.883 0.183 39.095 49.596 50.279
DDFM [49] 0.158 43.119 44.095 6.897 0.172 44.490 44.545 37.452 0.186 34.059 32.807 6.029 0.209 48.479 32.377 6.399
DRMF [32] 0.171 45.524 43.693 6.230 0.208 43.982 45.561 40.315 0.192 44.617 44.085 5.744 0.187 44.714 43.650 6.354
EMMA [50] 0.169 39.092 46.046 6.517 0.154 48.724 43.113 53.861 0.172 39.666 44.466 6.184 0.153 42.802 44.239 6.382
LRRNet [12] 0.170 48.571 48.973 7.363 0.151 48.718 45.266 51.605 0.158 48.274 37.090 7.295 0.137 46.511 36.610 7.702
SegMiF [16] 0.147 46.139 45.019 7.281 0.160 44.659 51.427 44.448 0.180 42.664 39.496 6.669 0.169 49.887 39.247 6.855
Text-IF [38] 0.178 50.568 50.271 6.956 0.187 49.299 49.266 47.032 0.169 46.647 48.491 6.256 0.161 49.019 47.755 6.085
Text-Difus[41] 0.175 52.788 53.073 7.470 0.165 50.092 51.715 39.429 0.203 49.278 49.479 6.982 0.194 51.762 49.288 7.089
VI (OE) and IR (LC) VI(Low-light and Noise, LN) VI (RH) and IR (RN) VI (LL) and IR (SN)
DDFM [49] 0.168 43.814 41.894 36.095 0.172 48.293 31.791 6.298 0.151 33.440 32.134 37.342 0.189 36.433 42.630 5.776
DRMF [32] 0.184 42.399 39.374 40.847 0.201 44.363 43.063 5.875 0.174 43.663 43.858 37.997 0.142 38.241 41.049 5.280
EMMA [50] 0.130 39.892 42.076 43.362 0.174 42.201 43.382 5.838 0.165 39.146 44.458 51.205 0.130 37.367 43.888 6.318
LRRNet [12] 0.136 47.209 42.636 46.684 0.144 46.386 35.779 7.306 0.128 47.954 36.831 49.917 0.154 38.426 35.970 7.007
SegMiF [16] 0.114 44.021 42.256 33.647 0.136 49.178 38.570 5.819 0.147 42.354 39.156 31.717 0.151 41.287 37.079 6.767
Text-IF [38] 0.174 48.808 47.998 48.848 0.217 48.100 47.510 5.204 0.158 45.821 47.626 46.543 0.140 41.429 46.220 5.525
Text-Difuse [41] 0.131 49.021 50.980 47.640 0.185 50.775 48.610 6.440 0.181 48.645 48.937 38.808 0.161 47.734 48.448 6.738

### v3

# ControlFusion: A Controllable Image Fusion Network with Language-Vision Degradation Prompts

@@ Abstract
Current image fusion methods struggle with real-world composite degradations and lack the flexibility to accommodate user-specific needs. To address this, we propose ControlFusion, a controllable fusion network guided by language-vision prompts that adaptively mitigates composite degradations. On the one hand, we construct a degraded imaging model based on physical mechanisms, such as the Retinex theory and atmospheric scattering principle, to simulate composite degradations and provide a data foundation for addressing realistic degradations. On the other hand, we devise a prompt-modulated restoration and fusion network that dynamically enhances features according to degradation prompts, enabling adaptability to varying degradation levels. To support user-specific preferences in visual quality, a text encoder is incorporated to embed user-defined degradation types and levels as degradation prompts. Moreover, a spatial-frequency collaborative visual adapter is designed to autonomously perceive degradations from source images, thereby reducing complete reliance on user instructions. Extensive experiments demonstrate that ControlFusion outperforms SOTA fusion methods in fusion quality and degradation handling, particularly under real-world and compound degradations. The source code is publicly available at https://github.com/Linfeng-Tang/ControlFusion.
footnotetext:  Equal contribution;  * Corresponding author.
@@ 1 Introduction
Image fusion is a crucial technique in image processing. By effectively leveraging complementary information from multiple sources, the limitations associated with information collection by single-modal sensors can be significantly mitigated [41]. Among the research areas in this field, infrared-visible image fusion (IVIF) has garnered considerable attention. IVIF integrates essential thermal information from infrared (IR) images with the intricate texture details from visible (VI) images, offering a richer and more complete depiction of the scene [46]. By harmonizing diverse information and producing visually striking results, IVIF has found extensive applications in areas such as security surveillance [47], military detection [22], scene understanding [44], and assisted driving [1]etc.
@@
Recently, IVIF has emerged as a focal point of research, leading to rapid advancements in related methods. Based on the adopted network architectures, these methods can be categorized into convolutional neural network-based [29, 51], autoencoder-based [10, 11], generative adversarial network-based [19, 15], Transformer-based [20, 45], and diffusion model-based [50, 32] methods. In addition, from a functional perspective, they can be grouped into visual-oriented [29, 51], joint registration-fusion [27, 37], semantic-driven [28, 16], and degradation-robust [31, 42, 43] schemes.
Figure 1: Comparison across real-world, composite degradation, and varying degradation levels.
@@
Notably, although degradation-robust schemes can mitigate degradation to some extent, certain limitations exist. First, existing restoration-fusion methods often adopt simplistic strategies for training data construction, overlooking the domain gap between simulated data and realistic images, which hampers their generalizability in practical scenarios, as illustrated in Fig. 1 (I). Second, they are tailored for specific or single types of degradation, making them ineffective in handling more complex composite degradations, as illustrated in Fig. 1 (II). Finally, as shown in Fig. 1 (III), existing methods lack degradation level modeling, causing a sharp decline in performance as degradation intensifies. Moreover, they lack flexibility to adapt fusion results to diverse user preferences.
@@
To overcome these limitations, we propose a versatile and controllable fusion model based on language-vision prompts, termed ControlFusion. On the one hand, we introduced a physics-driven degradation imaging model that differs from existing approaches by simultaneously simulating degradation processes for both visible and infrared modalities with high precision. This model not only effectively narrows the gap between synthetic data and real-world images but also provides crucial data support for addressing complex multi-modal composite degradation challenges. On the other hand, we develop a prompt-modulated image restoration and fusion network to generate high-quality fusion results. The prompt-modulated module enables our network to dynamically adjust feature distribution based on degradation characteristics (which can be specified by users), achieving robust feature enhancement. This also allows our method to respond to diverse user requirements. Furthermore, we devise a spatial-frequency visual adapter that combines frequency-domain degradation priors to directly extract degradation cues from degraded inputs, eliminating the heavy reliance on user instructions, as customizing prompts for each scene is time-consuming and labor-intensive. As shown in Fig. 1, benefiting from the aforementioned designs, our method excels in handling both real-world and composite degradation scenarios and can flexibly respond to user needs. In summary, our main contributions are as follows:
@@
$\bullet$ We propose ControlFusion, a versatile image restoration and fusion framework that uniformly models diverse degradation types and degrees, using textual and visual prompts as a medium. Specifically, its controllability enables it to respond to user-specific customization needs.
@@
$\bullet$ A spatial-frequency visual adapter is devised to integrate frequency characteristics and directly extract text-aligned degradation prompts from visual images, enabling automated deployment.
@@
$\bullet$ A physics-driven imaging model is developed, integrating physical mechanisms such as the Retinex theory and atmospheric scattering principle to bridge the gap between synthetic data and real-world images, while taking into account the degradation simulation of infrared-visible dual modalities.
@@ 2 Related Work
Image Fusion. Earlier visual-oriented fusion approaches primarily concentrated on merging complementary information from multiple modalities and improving visual fidelity. The mainstream network architectures primarily include CNN-based [14, 48], AE-based [11, 49], GAN-based [19, 15], Transformers [20] and diffusion models [50]. Furthermore, several schemes including joint registration and fusion [33, 37], semantic-driven [28, 15], and degradation-robust [39, 42, 38] methods, are proposed to broaden the practical applications of image fusion. For instance, Tang et al. [30] and Liu et al. [17] developed corresponding solutions for illumination distortions and noise interference, respectively. In addition, [39] proposed a image restoration and fusion network to address various degradations. However, it fails to handle mixed degradations and struggles to generalize to real-world scenarios. In particular, it relies on tedious manual efforts to customize text prompts for each scene, hindering large-scale automated deployment.
Image Restoration with Vision-Language Models. With the advancement of deep learning toward multimodal integration, the image restoration field has progressively evolved into a new paradigm of text-driven schemes. CLIP [25] establishes visual-textual semantic representation alignment through dual-stream Transformer encoders pre-trained on 400 million image-text pairs. This framework lays the foundation for the widespread adoption of Prompt Engineering in computer vision. Recent studies integrate CLIP encoders with textual user instructions, proposing generalized restoration frameworks including PromptIR [23], AutoDIR [8], and InstructIR [2], which effectively handle diverse degradation types. Moreover, SPIRE [24] further introduces fine-grained textual restoration cues by quantifying degradation severity levels and supplementing semantic information for precision restoration. To eliminate reliance on manual guidance, Luo et al. [18] developed DA-CLIP by fine-tuning CLIP on mixed-degradation datasets, enabling degradation-aware embeddings to facilitate distortion handling. However, existing unified restoration methods are primarily designed for natural images, which exhibit notable limitations when processing multimodal images (e.g., infrared-visible).
@@ 3 Physics-driven Degraded Imaging Model
Due to the differences in the imaging mechanisms of infrared and visible, the types of degradation they face also vary. To tackle the complex and variable degradation challenges, we propose a physics-driven imaging model for infrared and visible images aimed at reducing the domain gap between simulated data and real-world imagery. Infrared (IR) images usually suffer from sensor-related interference, such as stripe noise and low contrast. While Visible (VI) images are typically degraded by illumination conditions (low light, over-exposure), weather (rain, haze), and sensor-related issues (noise, blur). Given a clear image $I_{m}(m\in\{ir,vi\})$, the proposed imaging model can be mathematically represented as:
@@
Based on Eq. (1), we construct a multi-modal composite Degradation Dataset with four Levels (DDL-12), comprising 12 distinct degradations. We selecte $2,050$ high-quality clear images from RoadScene [36], LLVIP [7], and MSRS [29] datasets. Subsequently, the degraded imaging model is employed to synthesize degraded images with 12 types and 4 levels of degradation. Finally, the DDL-12 dataset contains approximately $48,000$ training image pairs and $4,800$ test image pairs.
Figure 2: The overall framework of our controllable image fusion network.
@@ 4.1 Problem Formulation
Given two source images $I_{ir}\in\mathbb{R}^{H\times W\times 1}$ and $I_{vi}\in\mathbb{R}^{H\times W\times 3}$, the typical fusion paradigm employs a network $\mathcal{N}_{f}$ to synthesize the fused image $I_{f}$, expressed as: $I_{f}=\mathcal{N}_{f}(I_{ir},I_{vi})$. However, in complex scenarios, source images usually suffer from degradation interference, thus the advanced fusion paradigm must take both image restoration and fusion into account. While the concatenation approach is straightforward, it does not model recovery and fusion as an end-to-end optimization process, yielding suboptimal outcomes. As illustrated in Fig. 2, we propose a controllable paradigm that couples restoration and fusion with degradation prompts, termed ControlFusion, to overcome this limitation. On the one hand, the coupled approach enhances task synergy. On the other hand, leveraging degradation prompts to modulate the restoration and fusion process enables the network to adapt to diverse degradation distributions while meeting user-specific customization requirements. The proposed restoration and fusion paradigm can be defined as: $I_{f}=\mathcal{N}_{rf}(I_{ir},I_{vi},p~|~\Omega)$, where $\mathcal{N}_{rf}$ is a restoration and fusion network, $P$ denotes degradation prompts, and $\Omega$ indicates the degradation set. Our ControlFusion employs a two-stage optimization protocol. Stage I aligns textual prompts with visual embedding, while Stage II optimizes the overall framework.
@@ 4.2 Stage I: Textual-Visual Prompts Alignment
Spatial-Frequency Collaborative Visual Adapter. To achieve unified multimodal degradation representation, we first employ text semantics for explicit degradation modeling. Using a pre-trained CLIP [25] text encoder $\mathcal{E}_{t}$, the user instruction $T_{instr}$ is converted into text embedding: $p_{text}=\mathcal{E}_{t}(T_{instr})$. However, text-dependent models limit deployment flexibility. We therefore develop a spatial-frequency collaborative visual adapter (SFVA) to extract visual embeddings directly from images while maintaining semantic alignment with $p_{text}$.
As shown in Fig. 3, degraded images exhibit distinct spectral characteristics, indicating frequency features contain rich degradation priors. Our SFVA contains dual branches: In the frequency branch, Fast Fourier Transform (FFT) and CNN extract frequency features:
$F_{fre}^{m}$ $=\textstyle{\sum_{x=0}^{W-1}\sum_{y=0}^{H-1}D_{m}(x,y)e^{-j2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)},}~F_{fre}=\text{Linear}\left(\text{Conv}\left([F_{fre}^{ir},F_{fre}^{vi}]\right)\right),$ (6)
where $\cdot$$\cdot$] denotes channel-wise concatenation. The spatial branch employs cropping/downsampling augmentation and CNN to extract spatial features $F_{spa}$.
The fused visual embedding $p_{vis}$ is obtained through concatenation and linear projection of $F_{fre}$ and $F_{spa}$. To ensure semantic consistency between visual/text embeddings, we apply MSE loss $\mathcal{L}_{mse}$ and cosine similarity loss $\mathcal{L}_{cos}$:
where $\lambda_{1}$ and $\lambda_{2}$ balance loss components. This design enables automatic degradation-aware adaptation while preserving text-aligned semantics.
@@ 4.3.1 Network Architectures
[…]
@@
$F_{f}^{ir}\!=\!\text{softmax}\Big(\!\frac{Q_{vi}K_{ir}}{\sqrt{d_{k}}}\!\Big)V_{ir},F_{f}^{vi}\!=\!\text{softmax}\Big(\!\frac{Q_{ir}K_{vi}}{\sqrt{d_{k}}}\!\Big)V_{vi},$ (9)
where $d_{k}$ is scaling factor. We concatenate $F_{f}^{ir}$ and $F_{f}^{vi}$ to get fusion features: $F_{f}=[F_{f}^{ir},F_{f}^{vi}]$.
[…]